OpenStudy (anonymous):

May someone help me study the surface areas, volumes, and change in dimensions for Cones, Cylinders and Spheres? Please?

OpenStudy (lgbasallote):

cone: \[SA = \pi r s + \pi r^2\] \[V = \frac 13 \pi r^2 h\] where r = radius s = slant height h = altitude Cylinder \[SA = 2 \pi r h + 2\pi r^2\] \[V = \pi r^2 h\] where: r = radius h = height Sphere: \[SA = 4\pi r ^2\] \[V = \frac 43 \pi r^3\] where: r = radius does that help?

OpenStudy (anonymous):

OK can I help @DarlingDare

OpenStudy (anonymous):

Thank you! :)